OpenStudy (anonymous):

Prove directly: Give n that a, b, c is a Pythagorean triple that qa, qb, and qc is a Pythagorean triple where q is an element of R and q does not equal 0.

OpenStudy (anonymous):

What happens when you enlarge any triangle by a factor of 2? (sides a,b,c) The sides become 2a, 2b and 2c. Same here. Except the Pythagorean triple implies a right triangle. And the factor is q. Starting with a right triangle with sides a,b,c we enlarge by a factor of q. The sides become qa, qb and qc. And what about the triangle? Still a right triangle. Therefore qa, qb and qc are a Pythagorean triple. Hope that makes sense :)

jhonyy9 (jhonyy9):

for an example 3,4 and 5 what mean that 9+16=25 so than q=2 and will get 2*3=6 ,2*4=8 and 2*5=10 what mean that 36+64=100 hope now this is more clearly !!!

OpenStudy (anonymous):

Didn't I solve it for you two hours ago?